I have a problem where I have to find the molar mass of something but I don't know where to start, this is everything I have and some overview on the instructions.
Equivalence point pH 8.79
Half point pH 4.33
pKa=4.33
Ka= 4.68E-5
Initial pH of acid 2.69
1.505 g of unknown acid was dissolved in a sufficient amount of water to dissolve the solid, and pitted 20mL into a beaker and stared titrating with NaOH 0.09695 M
Answer #1

Initial pH = 1/2 (pKa - Log[acid])

i.e. 2.69 = 1/2 (4.33 - Log[acid])

i.e. Log[acid] = -1.05

i.e. [acid] = 0.089 M

i.e. no. of moles of acid/volume of solution = 0.089 mol/L

i.e. (1.505 g/MW of acid)*1000/20 = 0.089

i.e. The MW of the unknown acid = 844.32 g/mol